Erol invests R8 350 in an account that pays simple interest. After six years, the amount that he receives back is R12 859. The simple interest rate on the investment, rounded to two decimal places, is

1.     0,75% per year.

2.     45,09% per year.

3.     1,08% per year.

4.     9,00% per year.

5.     none of the above.


1
Expert's answer
2021-05-17T15:01:27-0400

"12859 -8350 = 8350 \times r \times 6 \\

4509 = 50100r \\

r = \frac{4509}{50100} = 0.09 = 9 \; \%"

Answer: option 4. 9,00% per year.